OpenStudy (anonymous):

what is the vertex formula of x^2-10x+12 and the turning point...

OpenStudy (anonymous):

god knows

OpenStudy (anonymous):

I'm sure I would know the answer if you were god..........

OpenStudy (anonymous):

magic formula is \[-\frac{b}{2a}\]

OpenStudy (anonymous):

in this case a = 1, b = -10 and \[-\frac{b}{2a}=5\]

OpenStudy (anonymous):

is first coordinate. second is \[5^2-10\times 5+12=25-50+12=-25+12=-13\]

OpenStudy (anonymous):

so that is the turning point as well

OpenStudy (anonymous):

whats whats the vertex and whats the turning point?!

OpenStudy (anonymous):

?

OpenStudy (anonymous):

the point is (5,-13) that is where the parabola sits. so is also the turning point

OpenStudy (anonymous):

Well I need the vertex formula.